# EMMA-500: Enhancing Massively Multilingual Adaptation of Large Language Models

### v1

@@ D.4 Math
Llama 2 7B 0.0636 0.0200 0.0089 0.0760 0.0168 0.1600 0.0232 0.1240 0.0209 0.0920 0.0183 0.0360 0.0118 0.0800 0.0172 0.0200 0.0089 0.0080 0.0056 0.0160 0.0080 0.0800 0.0172
Llama 2 7B Chat 0.1091 0.0240 0.0097 0.1760 0.0241 0.2720 0.0282 0.1920 0.0250 0.1880 0.0248 0.0400 0.0124 0.1520 0.0228 0.0200 0.0089 0.0080 0.0056 0.0280 0.0105 0.1160 0.0203
CodeLlama 2 7B 0.0664 0.0160 0.0080 0.0920 0.0183 0.1320 0.0215 0.1080 0.0197 0.1080 0.0197 0.0440 0.0130 0.0760 0.0168 0.0120 0.0069 0.0160 0.0080 0.0600 0.0151 0.0400 0.0124
LLaMAX Llama 2 7B 0.0362 0.0360 0.0118 0.0360 0.0118 0.0720 0.0164 0.0360 0.0118 0.0520 0.0141 0.0280 0.0105 0.0280 0.0105 0.0080 0.0056 0.0080 0.0056 0.0200 0.0089 0.0480 0.0135
LLaMAX Llama 2 7B Alpaca 0.0635 0.0320 0.0112 0.0480 0.0135 0.1520 0.0228 0.0960 0.0187 0.0520 0.0141 0.0440 0.0130 0.0360 0.0118 0.0480 0.0135 0.0040 0.0040 0.0480 0.0135 0.0680 0.0160
MaLA-500 Llama 2 10B v1 0.0073 0.0000 0.0000 0.0040 0.0040 0.0040 0.0040 0.0040 0.0040 0.0160 0.0080 0.0160 0.0080 0.0240 0.0097 0.0040 0.0040 0.0000 0.0000 0.0000 0.0000 0.0080 0.0056
MaLA-500 Llama 2 10B v2 0.0073 0.0000 0.0000 0.0040 0.0040 0.0040 0.0040 0.0040 0.0040 0.0160 0.0080 0.0160 0.0080 0.0240 0.0097 0.0040 0.0040 0.0000 0.0000 0.0000 0.0000 0.0080 0.0056
YaYi Llama 2 7B 0.0722 0.0280 0.0105 0.0840 0.0176 0.1680 0.0237 0.1240 0.0209 0.1040 0.0193 0.0480 0.0135 0.0720 0.0164 0.0200 0.0089 0.0120 0.0069 0.0320 0.0112 0.1240 0.0209
TowerBase Llama 2 7B 0.0616 0.0360 0.0118 0.0800 0.0172 0.1120 0.0200 0.0840 0.0176 0.0760 0.0168 0.0360 0.0118 0.0840 0.0176 0.0120 0.0069 0.0080 0.0056 0.0280 0.0105 0.0920 0.0183
TowerInstruct Llama 2 7B 0.0824 0.0120 0.0069 0.1280 0.0212 0.1880 0.0248 0.1520 0.0228 0.1200 0.0206 0.0240 0.0097 0.1360 0.0217 0.0160 0.0080 0.0120 0.0069 0.0320 0.0112 0.1080 0.0197
Occiglot Mistral 7B v0.1 0.1407 0.0240 0.0097 0.2160 0.0261 0.3320 0.0298 0.2680 0.0281 0.2240 0.0264 0.0640 0.0155 0.1720 0.0239 0.0320 0.0112 0.0120 0.0069 0.0600 0.0151 0.1120 0.0200
Occiglot Mistral 7B v0.1 Instruct 0.2216 0.0400 0.0124 0.3320 0.0298 0.4320 0.0314 0.4200 0.0313 0.3160 0.0295 0.1680 0.0237 0.2440 0.0272 0.0560 0.0146 0.0160 0.0080 0.0840 0.0176 0.2480 0.0274
BLOOM 7B 0.0229 0.0200 0.0089 0.0160 0.0080 0.0480 0.0135 0.0240 0.0097 0.0240 0.0097 0.0040 0.0040 0.0440 0.0130 0.0160 0.0080 0.0200 0.0089 0.0200 0.0089 0.0240 0.0097
BLOOMZ 7B 0.0215 0.0200 0.0089 0.0200 0.0089 0.0240 0.0097 0.0320 0.0112 0.0320 0.0112 0.0200 0.0089 0.0160 0.0080 0.0240 0.0097 0.0200 0.0089 0.0160 0.0080 0.0120 0.0069
mGPT 0.0142 0.0000 0.0000 0.0200 0.0089 0.0200 0.0089 0.0160 0.0080 0.0240 0.0097 0.0120 0.0069 0.0200 0.0089 0.0280 0.0105 0.0000 0.0000 0.0000 0.0000 0.0160 0.0080
mGPT 13B 0.0153 0.0000 0.0000 0.0200 0.0089 0.0360 0.0118 0.0160 0.0080 0.0240 0.0097 0.0120 0.0069 0.0160 0.0080 0.0080 0.0056 0.0040 0.0040 0.0080 0.0056 0.0200 0.0089
XGLM 7.5B 0.0116 0.0000 0.0000 0.0120 0.0069 0.0000 0.0000 0.0080 0.0056 0.0000 0.0000 0.0040 0.0040 0.0160 0.0080 0.0200 0.0089 0.0240 0.0097 0.0160 0.0080 0.0280 0.0105
YaYi 7B 0.0302 0.0480 0.0135 0.0320 0.0112 0.0480 0.0135 0.0360 0.0118 0.0400 0.0124 0.0200 0.0089 0.0120 0.0069 0.0160 0.0080 0.0280 0.0105 0.0080 0.0056 0.0600 0.0151
Llama 3 8B 0.2813 0.1920 0.0250 0.3840 0.0308 0.5400 0.0316 0.4600 0.0316 0.3600 0.0304 0.0200 0.0089 0.3800 0.0308 0.2520 0.0275 0.0640 0.0155 0.3920 0.0309 0.0400 0.0124
Llama 3.1 8B 0.2731 0.1920 0.0250 0.4200 0.0313 0.5280 0.0316 0.4280 0.0314 0.3480 0.0302 0.0280 0.0105 0.3720 0.0306 0.2520 0.0275 0.0560 0.0146 0.3560 0.0303 0.0160 0.0080
Gemma 7B 0.3578 0.3640 0.0305 0.4080 0.0311 0.6080 0.0309 0.4520 0.0315 0.4200 0.0313 0.0360 0.0118 0.3920 0.0309 0.3880 0.0309 0.2480 0.0274 0.4400 0.0315 0.0480 0.0135
Gemma 2 9B 0.4469 0.4880 0.0317 0.5360 0.0316 0.7200 0.0285 0.6040 0.0310 0.5000 0.0317 0.0200 0.0089 0.5000 0.0317 0.5360 0.0316 0.4440 0.0315 0.5520 0.0315 0.0160 0.0080
Qwen 2 7B 0.5147 0.4520 0.0315 0.6600 0.0300 0.8200 0.0243 0.7520 0.0274 0.7160 0.0286 0.0280 0.0105 0.6840 0.0295 0.1720 0.0239 0.1760 0.0241 0.5840 0.0312 0.6080 0.0309
Qwen 1.5 7B 0.3036 0.1280 0.0212 0.4320 0.0314 0.5840 0.0312 0.5040 0.0317 0.4200 0.0313 0.1120 0.0200 0.3680 0.0306 0.0880 0.0180 0.0400 0.0124 0.2560 0.0277 0.2600 0.0278
EMMA-500 Llama 2 7B 0.1809 0.0840 0.0176 0.2240 0.0264 0.3760 0.0307 0.2560 0.0277 0.2160 0.0261 0.0800 0.0172 0.2280 0.0266 0.2120 0.0259 0.0240 0.0097 0.1160 0.0203 0.1640 0.0235

### v2

@@ Appendix E Detailed Results
Llama 2 7B 6.36 2.00 0.89 7.60 1.68 16.00 2.32 12.40 2.09 9.20 1.83 3.60 1.18 8.00 1.72 2.00 0.89 0.80 0.56 1.60 0.80 8.00 1.72
Llama 2 7B Chat 10.91 2.40 0.97 17.60 2.41 27.20 2.82 19.20 2.50 18.80 2.48 4.00 1.24 15.20 2.28 2.00 0.89 0.80 0.56 2.80 1.05 11.60 2.03
CodeLlama 2 7B 6.64 1.60 0.80 9.20 1.83 13.20 2.15 10.80 1.97 10.80 1.97 4.40 1.30 7.60 1.68 1.20 0.69 1.60 0.80 6.00 1.51 4.00 1.24
LLaMAX Llama 2 7B 3.62 3.60 1.18 3.60 1.18 7.20 1.64 3.60 1.18 5.20 1.41 2.80 1.05 2.80 1.05 0.80 0.56 0.80 0.56 2.00 0.89 4.80 1.35
LLaMAX Llama 2 7B Alpaca 6.35 3.20 1.12 4.80 1.35 15.20 2.28 9.60 1.87 5.20 1.41 4.40 1.30 3.60 1.18 4.80 1.35 0.40 0.40 4.80 1.35 6.80 1.60
MaLA-500 Llama 2 10B v1 0.73 0.00 0.00 0.40 0.40 0.40 0.40 0.40 0.40 1.60 0.80 1.60 0.80 2.40 0.97 0.40 0.40 0.00 0.00 0.00 0.00 0.80 0.56
MaLA-500 Llama 2 10B v2 0.73 0.00 0.00 0.40 0.40 0.40 0.40 0.40 0.40 1.60 0.80 1.60 0.80 2.40 0.97 0.40 0.40 0.00 0.00 0.00 0.00 0.80 0.56
YaYi Llama 2 7B 7.22 2.80 1.05 8.40 1.76 16.80 2.37 12.40 2.09 10.40 1.93 4.80 1.35 7.20 1.64 2.00 0.89 1.20 0.69 3.20 1.12 12.40 2.09
TowerBase Llama 2 7B 6.16 3.60 1.18 8.00 1.72 11.20 2.00 8.40 1.76 7.60 1.68 3.60 1.18 8.40 1.76 1.20 0.69 0.80 0.56 2.80 1.05 9.20 1.83
TowerInstruct Llama 2 7B 8.24 1.20 0.69 12.80 2.12 18.80 2.48 15.20 2.28 12.00 2.06 2.40 0.97 13.60 2.17 1.60 0.80 1.20 0.69 3.20 1.12 10.80 1.97
Occiglot Mistral 7B v0.1 14.07 2.40 0.97 21.60 2.61 33.20 2.98 26.80 2.81 22.40 2.64 6.40 1.55 17.20 2.39 3.20 1.12 1.20 0.69 6.00 1.51 11.20 2.00
Occiglot Mistral 7B v0.1 Instruct 22.16 4.00 1.24 33.20 2.98 43.20 3.14 42.00 3.13 31.60 2.95 16.80 2.37 24.40 2.72 5.60 1.46 1.60 0.80 8.40 1.76 24.80 2.74
BLOOM 7B 2.29 2.00 0.89 1.60 0.80 4.80 1.35 2.40 0.97 2.40 0.97 0.40 0.40 4.40 1.30 1.60 0.80 2.00 0.89 2.00 0.89 2.40 0.97
BLOOMZ 7B 2.15 2.00 0.89 2.00 0.89 2.40 0.97 3.20 1.12 3.20 1.12 2.00 0.89 1.60 0.80 2.40 0.97 2.00 0.89 1.60 0.80 1.20 0.69
mGPT 1.42 0.00 0.00 2.00 0.89 2.00 0.89 1.60 0.80 2.40 0.97 1.20 0.69 2.00 0.89 2.80 1.05 0.00 0.00 0.00 0.00 1.60 0.80
mGPT 13B 1.53 0.00 0.00 2.00 0.89 3.60 1.18 1.60 0.80 2.40 0.97 1.20 0.69 1.60 0.80 0.80 0.56 0.40 0.40 0.80 0.56 2.00 0.89
YaYi 7B 3.02 4.80 1.35 3.20 1.12 4.80 1.35 3.60 1.18 4.00 1.24 2.00 0.89 1.20 0.69 1.60 0.80 2.80 1.05 0.80 0.56 6.00 1.51
Aya 23 8B 24.71 6.00 1.51 40.80 3.11 48.80 3.17 41.20 3.12 40.00 3.10 28.40 2.86 37.20 3.06 5.60 1.46 2.40 0.97 8.40 1.76 6.80 1.60
Aya expanse 8B 41.45 21.20 2.59 68.80 2.94 78.80 2.59 76.80 2.68 65.20 3.02 29.60 2.89 67.60 2.97 11.20 2.00 2.40 0.97 20.00 2.53 2.80 1.05
Llama 3 8B 28.13 19.20 2.50 38.40 3.08 54.00 3.16 46.00 3.16 36.00 3.04 2.00 0.89 38.00 3.08 25.20 2.75 6.40 1.55 39.20 3.09 4.00 1.24
Llama 3.1 8B 27.31 19.20 2.50 42.00 3.13 52.80 3.16 42.80 3.14 34.80 3.02 2.80 1.05 37.20 3.06 25.20 2.75 5.60 1.46 35.60 3.03 1.60 0.80
Gemma 7B 35.78 36.40 3.05 40.80 3.11 60.80 3.09 45.20 3.15 42.00 3.13 3.60 1.18 39.20 3.09 38.80 3.09 24.80 2.74 44.00 3.15 4.80 1.35
Gemma 2 9B 44.69 48.80 3.17 53.60 3.16 72.00 2.85 60.40 3.10 50.00 3.17 2.00 0.89 50.00 3.17 53.60 3.16 44.40 3.15 55.20 3.15 1.60 0.80
Qwen 2 7B 51.47 45.20 3.15 66.00 3.00 82.00 2.43 75.20 2.74 71.60 2.86 2.80 1.05 68.40 2.95 17.20 2.39 17.60 2.41 58.40 3.12 60.80 3.09
Qwen 1.5 7B 30.36 12.80 2.12 43.20 3.14 58.40 3.12 50.40 3.17 42.00 3.13 11.20 2.00 36.80 3.06 8.80 1.80 4.00 1.24 25.60 2.77 26.00 2.78
EMMA-500 Llama 2 7B 18.09 8.40 1.76 22.40 2.64 37.60 3.07 25.60 2.77 21.60 2.61 8.00 1.72 22.80 2.66 21.20 2.59 2.40 0.97 11.60 2.03 16.40 2.35
Appendix F Open-Ended Generation
We have moved the open-ended generation task from the main text to the appendix due to the challenges in establishing a reliable evaluation method. Automated metrics, while useful, often fail to capture the full quality of open-ended responses, particularly in a multilingual and low-resource setting. Similarly, LLM-as-judge approaches can inherit biases from the models they rely on, leading to inconsistent or skewed assessments. Reward modeling, though promising, remains an active research area and is not yet robust enough for evaluating diverse multilingual outputs. Given these limitations, we opted to present the results of automatic metrics from the first version of this paper in the appendix for transparency while keeping the focus of the main text on tasks with more reliable evaluation methods
F.1 Evaluation Benchmarks
To facilitate the evaluation on open-ended generation. We build a benchmark, called PolyWrite, for writing tasks. The PolyWrite dataset has 51 writing tasks with the number of prompts per task shown in Figure 6. We use Google Translate to translate the English prompts into 240 languages and back-translate for translation quality assessment.
Figure 5: Writing tasks in the PolyWrite dataset.
Figure 6: Mean BLEU scores per language in the PolyWrite dataset.
F.2 Evaluation Results
Aya Evaluation
PolyWrite
This is a novel multilingual benchmark composed in this work for evaluating open-ended generation in 240 languages. We use ChatGPT to generate different prompts in English and use Google Translate to translate them into different languages for models to generate creative content. This benchmark consists of 31 writing tasks, such as storytelling and email writing, and 155 prompts in total. We back-translate the multilingual prompts to English, calculate the BLEU scores between original English prompts and back-translation, and filter out translated prompts with BLEU scores below 20, and the entire dataset contains a total of 35,751 prompts. We release the PolyWrite dataset on Huggingface [^48] https://huggingface.co/datasets/MaLA-LM/PolyWrite. The details of PolyWrite are described in Section F.1.
Table 27: Results on Aya (BLEU/Self-BLEU) and PolyWrite (Self-BLEU). EMMA-500 Llama 2 7B has higher average BLEU scores than all baselines on Aya.
Model Aya PolyWrite
Avg High Med-High Medium Med-Low Low Avg High Med-High Medium Med-Low Low
Llama 2 7B 1.24/0.74 1.27/0.57 1.47/0.60 0.86/0.37 1.17/0.45 0.77/1.87 0.5358 0.4282 0.6545 0.3769 0.4766 0.6587
Llama 2 7B Chat 1.17/1.29 1.46/1.15 1.36/1.15 0.69/1.03 1.14/1.14 0.54/2.23 1.1550 0.8640 0.8902 1.1877 1.4435 1.4167
CodeLlama 2 7B 1.22/1.21 1.31/1.19 1.40/1.00 0.85/0.84 1.23/0.56 0.78/2.57 1.0313 1.2052 1.2883 0.9191 0.9092 0.6798
LLaMAX Llama 2 7B 1.72/1.70 1.80/1.37 2.03/1.48 1.30/1.24 1.65/0.89 0.98/3.74 0.9564 1.0066 1.1655 0.8786 0.8363 0.7709
LLaMAX Llama 2 7B Alpaca 1.68/1.67 1.82/1.22 1.96/1.42 1.28/1.14 1.55/1.04 1.00/3.94 0.8086 0.7321 0.9517 0.8322 0.8351 0.4981
MaLA-500 Llama 2 10B v1 0.40/2.29 0.42/2.53 0.49/2.79 0.32/1.22 0.31/2.42 0.18/1.16 3.7079 3.7902 3.5066 2.4541 4.8052 3.5163
MaLA-500 Llama 2 10B v2 0.41/2.31 0.42/2.01 0.50/2.65 0.32/1.42 0.33/1.02 0.18/3.09 4.0059 3.1737 4.0148 3.0476 5.0652 3.8916
Yayi Llama 2 7B 1.65/0.61 1.84/0.82 1.88/0.62 1.26/0.62 1.53/0.45 1.02/0.41 0.6274 0.6207 0.6921 0.4169 0.6813 0.6352
TowerBase Llama 2 7B 1.44/0.83 1.45/0.64 1.67/0.56 1.19/0.49 1.46/0.38 0.88/2.54 0.4938 0.7268 0.4736 0.3945 0.4417 0.5396
TowerInstruct Llama 2 7B 1.55/0.93 1.80/0.85 1.82/0.78 1.15/0.65 1.21/0.54 0.85/1.97 0.7124 0.9565 0.7651 0.6492 0.5998 0.6615
Occiglot Mistral 7B v0.1 1.53/2.43 1.57/0.91 1.78/2.89 1.17/1.73 1.61/1.31 0.95/4.27 0.9647 0.8818 0.7975 0.9543 0.9563 1.4157
Occiglot Mistral 7B v0.1 Instruct 0.75/2.81 0.82/2.38 0.89/3.10 0.43/2.72 0.79/1.17 0.45/3.50 3.9033 5.3884 4.7140 3.7555 2.8444 2.9568
BLOOM 7B 0.85/1.17 0.92/1.20 0.96/1.32 0.73/1.05 0.78/1.27 0.55/0.72 1.3892 1.2845 1.6705 1.9685 1.1513 0.6600
BLOOMZ 7B 0.12/0.61 0.07/0.33 0.17/0.62 0.08/1.00 0.06/0.89 0.08/0.51 0.0024 0.0000 0.0005 0.0093 0.0000 0.0049
mGPT 1.24/0.55 1.22/0.64 1.47/0.59 0.91/0.48 1.21/0.60 0.84/0.29 0.7560 0.9222 0.6291 0.4534 0.8156 1.1134
mGPT-13B 1.42/0.57 1.42/0.80 1.63/0.58 1.00/0.48 1.53/0.44 1.03/0.36 0.7479 0.8483 0.7091 0.4230 0.7962 1.0201
Yayi 7B 1.05/0.39 1.22/0.38 1.18/0.41 0.76/0.42 1.01/0.54 0.67/0.22 0.5151 0.4266 0.3574 0.3283 0.6706 0.8574
Llama 3 8B 1.59/0.94 1.03/0.60 1.31/0.53 2.96/0.54 1.11/0.28 2.43/3.47 0.5796 0.5753 0.5921 0.7141 0.5586 0.4440
Llama 3.1 8B 1.85/1.08 1.41/0.95 1.60/0.64 3.11/0.52 1.33/0.39 2.52/3.52 0.7995 0.6805 0.8253 0.5044 0.6965 1.3585
Gemma 2 9B 1.55/0.82 1.59/0.94 1.73/0.93 1.38/0.70 1.33/0.47 1.21/0.65 1.1736 1.2347 1.2913 1.1616 0.9261 1.3240
Gemma 7B 1.29/0.57 1.41/0.62 1.39/0.66 1.16/0.40 1.26/0.50 0.93/0.40 1.0541 0.9629 1.1222 0.9275 1.1112 1.0284
Qwen 1.5 7B 1.93/1.13 1.66/0.85 1.64/0.65 3.14/0.79 1.51/0.58 2.45/3.69 0.6441 0.9398 0.7457 0.3797 0.4164 0.8738
Qwen 2 7B 1.99/1.13 1.84/0.94 1.69/0.70 3.29/0.72 1.36/0.44 2.47/3.53 0.5709 0.7763 0.6135 0.3695 0.4186 0.7989
EMMA-500 Llama 2 7B 2.93/1.54 2.90/1.29 2.75/0.83 3.79/0.82 2.86/0.95 2.87/5.09 0.9879 0.9833 0.7058 0.8465 1.3130 1.1702